Suppose a recent college graduate's first job allows her to deposit $200 at the end of each month in a savings plan that earns 6%, compounded monthly. This savings plan continues for 7 years before new obligations make it impossible to continue. If the accrued amount remains in the plan for the next 15 years without deposits or withdrawals, how much money will be in the account 22 years after the plan began? (Round your answer to the nearest cent.)
Answer #1

Rate per month =6%/12 =0.5%
PMT =200
Number of Periods =12*7 =84
FV =PMT*((1+r)^n-1)/r =200*((1+0.5%)^84-1)/0.5% =20814.7854

Number of Years from 7 to year 22 =15
Amount after in 22 years =Value in year 7*(1+0.5%)^(15*12) =20814.7854*(1+0.5%)^180 =51081.43